# The heat trace for domains with curved corners

### v2

@@ Abstract.
We derive the short-time expansion of the heat trace on planar curvilinear polygons through order $t^{1/2}$, with both Dirichlet and Neumann boundary conditions. For such domains we show that the coefficient of $t^{1/2}$ in the expansion splits into a boundary integral of $\kappa^{2}$ and a sum of local corner contributions, one for each vertex. Each curved corner contribution depends only on the interior angle $\alpha$ and on the limiting curvatures $\kappa_{\pm}$ on the adjacent sides. For Dirichlet boundary conditions, using a conformal model and a parametrix construction on the sector heat space, we express this contribution in the form $c_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$, where $c_{1/2}(\alpha)$ is given by the Hadamard finite part of an explicit trace over the exact sector. For right angles we compute $c_{1/2}(\pi/2)=1/(16\sqrt{\pi})$, which yields a closed formula for the $\sqrt{t}$ coefficient for domains with right-angled corners. Combined with previous work on the coefficients of $t^{-1}$, $t^{-1/2}$ and $t^{0}$ for such domains, this determines the remaining local heat trace coefficient for planar curvilinear polygons up to order $t^{1/2}$. As an inverse spectral application we extend a result in the literature by showing that any admissible curvilinear polygon that is Dirichlet isospectral to a polygon must itself be a polygon with straight sides.
@@ 1. Introduction
We can say substantially more about $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ whenever $\alpha\neq\pi$. Define
@@
(3) $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c_{1/2}(\alpha)r_{0}(\alpha,\kappa_{+},\kappa_{-})$
where $c_{1/2}(\alpha)$ is a function depending only on $\alpha$, whose general expression is given by a Hadamard renormalized trace on an exact sector:
@@
$\quad a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\,\rho}{\sqrt{1-\sigma}},\phi\right)\,b\left(\rho,\phi,\frac{R}{\sqrt{\sigma}},\theta\right)\,\,\rho\,d\rho\,d\phi\,\,R\,dR\,d\theta\,d\sigma.$
Here $a$ and $b$ are functions related to the exact heat kernel on a sector of angle $\alpha$, explicitly given by (43) and (58), and $\operatorname*{fp}$ denotes the Hadamard finite part at $R=\infty$ of the improper integral in $R$.
@@
(5) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
The factorization (3) shows that, away from $\alpha=\pi$, the corner contribution at order $t^{1/2}$ splits cleanly into an angular factor $c_{1/2}(\alpha)$ and a curvature factor $r_{0}(\alpha,\kappa_{+},\kappa_{-})$, so that all dependence on the limiting curvatures is encoded in a single scalar quantity. This theorem is not just an integral identity, but a structural description of how corner curvature enters the expansion.
@@
  • Each of $a$ and $b$ is a function only of $\alpha$ and the variables of integration, see (43), (58). Thus (4) is indeed independent of $\kappa_{\pm}$.
@@
  • Theorem 2 does not work for $\alpha=\pi$ because the conformal model that we use breaks down in that case, as can be seen from the fact that $r_{0}$ is undefined when $\alpha=\pi$. So the case $\alpha=\pi$, where the boundary has an abrupt change in second derivative, remains open.
@@ 1.1. Inverse spectral applications
A curvilinear polygon $\Omega$ is admissible if, for each corner $P_{j}$ with opening angle $\alpha_{j}$, either
$c_{1/2}(\alpha_{j})>0\quad\text{or}\quad\kappa_{j,+}=\kappa_{j,-}=0.$
@@
(7) $\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}c_{1/2}(\alpha_{j})r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})$
with $r_{0}$ defined in (2). Note that $\int_{\partial\Omega}\kappa^{2}\,ds\geq 0$. Moreover $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\geq 0$ by definition, and (for $\alpha_{j}\neq\pi$) one has $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ if and only if $\kappa_{j,+}=\kappa_{j,-}=0$. By admissibility, whenever $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})>0$ we necessarily have $c_{1/2}(\alpha_{j})>0$. Hence every term in (7) is non-negative. Since the $t^{1/2}$ coefficient vanishes, each term must vanish: $\int_{\partial\Omega}\kappa^{2}\,ds=0$ and $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ for all $j$. Therefore $\kappa\equiv 0$ on each smooth edge and $\kappa_{j,+}=\kappa_{j,-}=0$ at every corner, so $\partial\Omega$ is piecewise linear and $\Omega$ is a polygon. ∎
@@
The admissibility hypothesis is used only to ensure that whenever a corner has nonzero curvature on at least one side, the corresponding coefficient satisfies $c_{1/2}(\alpha_{j})>0$. The alternative condition $\kappa_{j,+}=\kappa_{j,-}=0$ requires only vanishing one-sided curvature limits; it does not assume that the boundary is straight in a full neighborhood of $P_{j}$.
@@
Observe from (5) that any polygon where all non-straight corners are right angles is admissible.
@@
In our argument, admissibility is used only through the implication
$(\kappa_{j,+},\kappa_{j,-})\neq(0,0)\quad\Longrightarrow\quad c_{1/2}(\alpha_{j})>0$
at each corner $P_{j}$. Any structural information on the sign of $\alpha\mapsto c_{1/2}(\alpha)$ would therefore allow one to weaken the definition of an admissible curvilinear polygon without changing the proofs.
We conjecture that $c_{1/2}(\alpha)>0$ for all $\alpha\in(0,\pi)$. In that case, every curvilinear polygon with corner angles in $(0,\pi)$ would be admissible in our sense.
@@ 1.2. Circular and annular sectors
(9) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
Note that if $\alpha_{j}=\pi/2$, then $\cos\alpha_{j}=0$ and $\csc\alpha_{j}=1$, and (2) reduces to
(10) $r_{0}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{2}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
Combining (9) and (10) gives
(11) $\mathcal{C}_{1/2}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{32\sqrt{\pi}}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
@@
Using (16),
(18) $a_{1/2}\bigl(A(r,R,\alpha)\bigr)=\frac{\alpha+16}{256\sqrt{\pi}}\left(\frac{1}{R}+\frac{1}{r}\right).$
@@
$a_{1/2}(\Omega_{1})=\frac{16+\frac{1}{4}}{256\sqrt{\pi}}\left(1+\frac{1}{3}\right)=\frac{65}{768\sqrt{\pi}},\qquad a_{1/2}(\Omega_{2})=\frac{16+\frac{1}{6}}{256\sqrt{\pi}}\left(\frac{1}{2}+\frac{1}{4}\right)=\frac{97}{2048\sqrt{\pi}},$
@@
(20) $\bigl(256\sqrt{\pi}\,a_{1/2}(\Omega)\bigr)\,(s^{2}-d^{2})=64s+\frac{8A}{d}.$
For each fixed $d>0$, this is a quadratic equation in $s$ with one positive root and one negative root. Thus each admissible choice of $d$ determines a unique $s>d$, and hence a unique triple $(r,R,\alpha)$.
@@ 3.4. Computations with $G$
(54) $\partial_{t}G=-\frac{a_{0}}{4}T^{-1}D^{2}S+\frac{a_{0}}{4}T^{-1}(R\partial_{R}+R^{\prime}\partial_{R}^{\prime})(D^{2}S)+O(1).$
@@
(55) $\mathcal{L}G=-\frac{a_{0}}{4}T^{-1}(1-R\partial_{R}-R^{\prime}\partial_{R^{\prime}}-2\Delta_{R,\theta})(D^{2}S)+O(1).$
@@ 3.4.1. The little trace and the convolution
The last thing to do is to address the dependence of this expression on $\kappa_{\pm}$. By (58), the kernel $b$ does appear to depend on $\theta_{0}$ through factors such as $\cos(\theta+\theta_{0})$. However, integrating in the angular variables removes this dependence. Let $I(\theta_{0})$ denote the right-hand side of (4). From (58), the $\theta_{0}$ dependence enters only through shifted angles of the form $\theta+\theta_{0}$ and $\phi+\theta_{0}$, while the remaining angular dependence, coming from the $D^{2}$ term, occurs through differences such as $\theta-\phi$. Differentiating under the integral sign therefore yields
$\frac{d}{d\theta_{0}}I(\theta_{0})=\iint_{(0,\alpha)\times(0,\alpha)}(\partial_{\theta}+\partial_{\phi})(\cdots)\,d\theta\,d\phi,$
where $(\cdots)$ denotes the $\theta,\phi$ dependent integrand in (4). Integrating by parts in $\theta$ and $\phi$ reduces this to boundary terms at $\theta,\phi\in\{0,\alpha\}$, and these vanish for Dirichlet boundary conditions since the sector kernel $a$ in (43) contains the factors $\varphi_{j}(\theta)$ and $\varphi_{j}(\phi)$ with $\varphi_{j}(0)=\varphi_{j}(\alpha)=0$. Hence $\frac{d}{d\theta_{0}}I(\theta_{0})=0$, so $c_{1/2}(\alpha)=I(\theta_{0})$ is independent of $\theta_{0}$. This justifies writing $c_{1/2}(\alpha)$, rather than $c_{1/2}(\alpha,\kappa_{-},\kappa_{+})$, on the left-hand side of (4).
@@ 4.1. The Dirichlet case
Combined with (3), and the fact that $r_{0}(\frac{\pi}{2},1,0)=\frac{1}{2}$, this identifies $c_{1/2}(\pi/2)$ and confirms that right-angled curved corners contribute nontrivially at order $t^{1/2}$. It remains only to prove the Lemmas.
@@ 4.3. A Neumann right-angle corner contribution
$\mathcal{C}^{N}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c^{N}_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$
holds for $\alpha=\pi/2$, then since $r_{0}(\pi/2,1,0)=\frac{1}{2}$ by (10), one obtains
$c^{N}_{1/2}\Bigl(\frac{\pi}{2}\Bigr)=\frac{\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)}{r_{0}(\pi/2,1,0)}=\frac{3}{16\sqrt{\pi}}.$

### v4

@@ Abstract.
We derive the short-time expansion of the heat trace on planar curvilinear polygons through order $t^{1/2}$, with both Dirichlet and Neumann boundary conditions. For such domains we show that the coefficient of $t^{1/2}$ in the expansion splits into a boundary integral of $\kappa^{2}$ and a sum of local corner contributions, one for each vertex. Each curved corner contribution depends only on the interior angle $\alpha$ and on the limiting curvatures $\kappa_{\pm}$ on the adjacent sides. For Dirichlet boundary conditions, using a conformal model and a parametrix construction on the sector heat space, we express this contribution in the form
$\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c_{1/2}(\alpha)\,\frac{\kappa_{+}+\kappa_{-}}{4\sin(\alpha/2)},$
where $c_{1/2}(\alpha)$ is given by the Hadamard finite part of an explicit trace over the exact sector. For right angles we compute $c_{1/2}(\pi/2)=\sqrt{2}/(16\sqrt{\pi})$, which yields the closed formula
$\mathcal{C}_{1/2}(\pi/2,\kappa_{+},\kappa_{-})=\frac{\kappa_{+}+\kappa_{-}}{32\sqrt{\pi}}.$
Combined with previous work on the coefficients of $t^{-1}$, $t^{-1/2}$ and $t^{0}$ for such domains, this determines the remaining local heat trace coefficient for planar curvilinear polygons up to order $t^{1/2}$. As an inverse spectral application we extend a result in the literature by showing that any admissible curvilinear polygon that is Dirichlet isospectral to a polygon must itself be a polygon with straight sides.
@@ 1. Introduction
We can say substantially more about $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ whenever $\alpha\neq\pi$.
For later use in the conformal model, define
@@
(3) $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c_{1/2}(\alpha)\,\frac{\kappa_{+}+\kappa_{-}}{4\sin(\alpha/2)}.$
Here $c_{1/2}(\alpha)$ depends only on $\alpha$ and is given by the Hadamard renormalized trace on the exact sector
@@
$\quad a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\,\rho}{\sqrt{1-\sigma}},\phi\right)\,b_{\mathrm{bis}}\left(\rho,\phi,\frac{R}{\sqrt{\sigma}},\theta\right)\,\,\rho\,d\rho\,d\phi\,\,R\,dR\,d\theta\,d\sigma,$
where $a$ is given by (44), $b_{\mathrm{bis}}$ is obtained from (59) by setting
$\theta_{0}=\pi-\frac{\alpha}{2},$
and $\operatorname*{fp}$ denotes the Hadamard finite part at $R=\infty$ of the improper integral in $R$.
@@
(5) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{\sqrt{2}}{16\sqrt{\pi}}.$
The factorization (3) shows that, away from $\alpha=\pi$, the corner contribution at order $t^{1/2}$ splits into an angular factor $c_{1/2}(\alpha)$ and the bisector-even curvature combination $(\kappa_{+}+\kappa_{-})/(4\sin(\alpha/2))$. This theorem is not just an integral identity, but a structural description of how the one-sided curvatures enter the expansion.
@@
  • Each of $a$ and $b_{\mathrm{bis}}$ is a function only of $\alpha$ and the variables of integration, see (44) and (59). Thus (4) is indeed independent of $\kappa_{\pm}$.
@@
  • Theorem 2 does not work for $\alpha=\pi$ because the conformal model that we use breaks down in that case, as can be seen from the fact that the parameter $r_{0}$ in (2) is undefined when $\alpha=\pi$. So the case $\alpha=\pi$, where the boundary has an abrupt change in second derivative, remains open.
@@ 1.1. Inverse spectral applications
A curvilinear polygon $\Omega$ is admissible if, for each corner $P_{j}$ with opening angle $\alpha_{j}$,
$c_{1/2}(\alpha_{j})\,\bigl(\kappa_{j,+}+\kappa_{j,-}\bigr)\geq 0.$
Since $\sin(\alpha_{j}/2)>0$ for $0<\alpha_{j}<2\pi$, this is equivalent to requiring that each corner term in (3) be nonnegative.
@@
(7) $\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}c_{1/2}(\alpha_{j})\,\frac{\kappa_{j,+}+\kappa_{j,-}}{4\sin(\alpha_{j}/2)}.$
Note that $\int_{\partial\Omega}\kappa^{2}\,ds\geq 0$. By admissibility, every corner term in (7) is also nonnegative. Since the full $t^{1/2}$ coefficient vanishes, each term in (7) must vanish. In particular,
$\int_{\partial\Omega}\kappa^{2}\,ds=0.$
Hence $\kappa\equiv 0$ on each smooth edge, so the one-sided limits $\kappa_{j,+}$ and $\kappa_{j,-}$ also vanish at every corner. Therefore $\partial\Omega$ is piecewise linear and $\Omega$ is a polygon. ∎
@@
The admissibility hypothesis is used only to ensure that each curved-corner contribution in (7) is nonnegative. Any criterion implying
$c_{1/2}(\alpha_{j})\,\bigl(\kappa_{j,+}+\kappa_{j,-}\bigr)\geq 0$
for every corner $P_{j}$ would be sufficient for the same proof.
@@
Observe from (5) that any convex curvilinear polygon whose non-straight corners are right angles is admissible.
@@
In our argument, admissibility is used only through the sign condition
$c_{1/2}(\alpha_{j})\,\bigl(\kappa_{j,+}+\kappa_{j,-}\bigr)\geq 0$
at each corner $P_{j}$. In particular, for convex corners the positivity of $\alpha\mapsto c_{1/2}(\alpha)$ is enough to guarantee admissibility.
Conjecture (Positivity conjecture). For all $\alpha\in(0,\pi)$, $c_{1/2}(\alpha)>0$.
If the positivity conjecture holds, then every convex curvilinear polygon is admissible, and Proposition 2 applies without further hypothesis on the corners.
@@ 1.2. Circular and annular sectors
(9) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{\sqrt{2}}{16\sqrt{\pi}}.$
Since $\sin(\pi/4)=\sqrt{2}/2$, (3) reduces in this case to
(10) $\mathcal{C}_{1/2}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{\kappa_{+}+\kappa_{-}}{32\sqrt{\pi}}.$
@@
Using (15), together with the fact that the two outer corners have $(\kappa_{+},\kappa_{-})=(1/R,0)$ and the two inner corners have $(\kappa_{+},\kappa_{-})=(-1/r,0)$, we obtain from (10)
(17) $a_{1/2}\bigl(A(r,R,\alpha)\bigr)=\frac{1}{256\sqrt{\pi}}\left(\frac{\alpha+16}{R}+\frac{\alpha-16}{r}\right).$
@@
$a_{1/2}(\Omega_{1})=\frac{1}{256\sqrt{\pi}}\left(\frac{16+\frac{1}{4}}{3}+\left(\frac{1}{4}-16\right)\right)=-\frac{31}{768\sqrt{\pi}},\qquad a_{1/2}(\Omega_{2})=\frac{1}{256\sqrt{\pi}}\left(\frac{16+\frac{1}{6}}{4}+\frac{\frac{1}{6}-16}{2}\right)=-\frac{31}{2048\sqrt{\pi}},$
@@
(19) $\bigl(256\sqrt{\pi}\,a_{1/2}(\Omega)\bigr)\,(s^{2}-d^{2})=\frac{8A}{d}-64d.$
For each fixed $d>0$, this determines $s^{2}$, and hence determines the unique value $s>d$. Thus each admissible choice of $d$ determines a unique triple $(r,R,\alpha)$.
@@ 3.1. The conformal model domain
Equivalently,
(26) $\kappa_{+}=2r_{0}\sin\theta_{0},\qquad\kappa_{-}=-2r_{0}\sin(\alpha+\theta_{0}).$
Hence
(27) $r_{0}\cos\left(\theta_{0}+\frac{\alpha}{2}\right)=-\frac{\kappa_{+}+\kappa_{-}}{4\sin(\alpha/2)}.$
@@ 3.4. Computations with $G$
(55) $\partial_{t}G=-\frac{r_{0}}{4}T^{-1}D^{2}S+\frac{r_{0}}{4}T^{-1}(R\partial_{R}+R^{\prime}\partial_{R}^{\prime})(D^{2}S)+O(1).$
@@
(56) $\mathcal{L}G=-\frac{r_{0}}{4}T^{-1}(1-R\partial_{R}-R^{\prime}\partial_{R^{\prime}}-2\Delta_{R,\theta})(D^{2}S)+O(1).$
@@ 3.4.1. The little trace and the convolution
The last thing to do is to address the dependence of this expression on $\kappa_{\pm}$. Let
$I_{\alpha}(\theta_{0})$
denote the renormalized integral obtained from the right-hand side of (4) by replacing $b_{\mathrm{bis}}$ with the general function $b$ from (59). Then the corner contribution produced by the conformal model is
$\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=r_{0}\,I_{\alpha}(\theta_{0}).$
By (59), every occurrence of $\theta_{0}$ is linear and enters only through shifted angles such as $\theta+\theta_{0}$ and $\phi+\theta_{0}$. Expanding these shifted trigonometric functions shows that
$I_{\alpha}(\theta_{0})=A(\alpha)\cos\theta_{0}+B(\alpha)\sin\theta_{0}$
for some coefficients $A(\alpha)$ and $B(\alpha)$.
Now reflect the sector across its bisector. This swaps the two boundary rays, hence swaps $\kappa_{+}$ and $\kappa_{-}$, but it does not change the heat trace coefficient because the reflected corner is isometric to the original one. In terms of the parameter $\theta_{0}$, the reflection sends
$\theta_{0}\mapsto-\alpha-\theta_{0},$
which is the transformation that exchanges the two formulas in (26). Therefore
$I_{\alpha}(\theta_{0})=I_{\alpha}(-\alpha-\theta_{0}).$
Combining this symmetry with the linear form above gives
$I_{\alpha}(\theta_{0})=C(\alpha)\cos\left(\theta_{0}+\frac{\alpha}{2}\right)$
for some function $C(\alpha)$. We now define
$c_{1/2}(\alpha):=I_{\alpha}\left(\pi-\frac{\alpha}{2}\right),$
which is exactly the renormalized trace formula (4). Since
$I_{\alpha}\left(\pi-\frac{\alpha}{2}\right)=-C(\alpha),$
we obtain
$I_{\alpha}(\theta_{0})=-c_{1/2}(\alpha)\cos\left(\theta_{0}+\frac{\alpha}{2}\right).$
Finally, substituting (27) gives
$\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=r_{0}I_{\alpha}(\theta_{0})=c_{1/2}(\alpha)\,\frac{\kappa_{+}+\kappa_{-}}{4\sin(\alpha/2)},$
which is (3).
@@ 4.1. The Dirichlet case
Combined with (3), this identifies
$c_{1/2}\left(\frac{\pi}{2}\right)=\frac{\sqrt{2}}{16\sqrt{\pi}},$
and confirms that right-angled curved corners contribute nontrivially at order $t^{1/2}$. It remains only to prove the Lemmas.
@@ 4.3. A Neumann right-angle corner contribution
$\mathcal{C}^{N}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c^{N}_{1/2}(\alpha)\,\frac{\kappa_{+}+\kappa_{-}}{4\sin(\alpha/2)}$
holds for $\alpha=\pi/2$, then since $(\kappa_{+},\kappa_{-})=(1,0)$ in (91), one obtains
$c^{N}_{1/2}\Bigl(\frac{\pi}{2}\Bigr)=\frac{\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)}{1/(2\sqrt{2})}=\frac{3\sqrt{2}}{16\sqrt{\pi}}.$